# Geometric Neural Network based on Phase Space for BCI-EEG decoding

### v2

@@ 3.4 Datasets
It turns out that it is essential to employ different datasets to obtain reliable results, especially when evaluating different ML techniques in EEG decoding, in order to avoid the bias of a single data set [13].
@@
We employed standard pre-processing steps designed for all datasets described in [10]. We applied band-pass filtering with the overlap-add method between $[8-32]$ Hz [77, 78], artifact rejection, and electrode standardization, bringing every channel to a zero mean and a unit standard deviation. Additionally, we opted to use the complete epoch duration, although this duration varies among datasets, as shown in Table 1. The pre-processing steps just mentioned are the same for all the pipelines considered in our study.
We concentrate on a binary classification task, aiming to discern between imagined movements of the right and left hand using an intra-subject/within-session evaluation, which is based on a $5-$fold cross-validation is conducted in each session.
Additionally, we focus on investigating our algorithm’s robustness with fewer electrode settings. For this purpose, our channel selection process was guided by two constraints: ensuring neurophysiological relevance and maintaining consistent electrode selection across datasets. We selected electrodes positioned over the sensorimotor region, particularly in the central (C) and centro-parietal (CP) lines, which are actively involved in the tasks. This choice is strongly supported by Brain-Computer Interface (BCI) research [78, 79, 80, 81, 82, 83, 84]. For consistency, we used EEG montages with varying numbers of electrodes, from 3 to 128. We standardized our analysis by using the minimal overlapping electrodes common to all montages: C3, Cz, and C4.
@@ 3.5 Baseline Comparison
These methods were implemented and trained using MOABB version 1.0  [10]. For in-depth information about the deep learning hyper-parameters, please consult Table A2. In order to obtain a fair time comparison, all the following results are computed on the same hardware, a Dell C6420 dual-Xeon Cascade Lake SP Gold 6240 @ 2.60GHz.
@@ 4 Results
In the following section, we analyze the results produced using within-session (WS) evaluation. This method is based on a 5-fold cross-validation performed on each session independently. In particular, we report two separate analyses.
@@
The first analysis is dedicated to contrasting the performance of Phase-SPDNet , with a specific focus on the exploitation of the covariance feature, against the state-of-the-art techniques in DL applied in EEG. The subsequent analysis, reported in the Appendix, shifts the focus towards the exploration of coherence as a feature for classification. Despite the ongoing investigations into coherence, we deliberately separated this analysis due to its comparatively lesser performance against the benchmarks achieved with covariance features. Nevertheless, the augmentation procedure is a significative methodology to improve the performance of both feature sets.
@@
For the first analysis, we tested the performance of Phase-SPDNet against the baseline comparison defined in 3.5 across a broad spectrum of datasets. The results are listed in Table 2, while the detailed statistical analysis can be found in Figure 3.
Table 2 shows ROC AUC results for right-hand versus left-hand classification, where bold numbers represent the best score in each dataset, revealing a consistent trend wherein our algorithm Phase-SPDNet surpasses the DL state-of-the-art in the context of reduced datasets. This behavior is true for all datasets, with the notable exception of the Schirrmeister2017 dataset, where it is marginally surpassed by ShallowNet, DeepNet and DynSpat-ShallowNet. It is important to note that the results of the first two, however, are not statistically significant for this dataset, as presented in Figure 3 (c), (d). The results obtained in this dataset by DynSpat-ShallowNet is statistically superior, as shown in Figure 3 (e). However, we are interested in computing a statistic behavior considering all datasets under consideration. As the number of subjects per dataset is actually too small to reach a decent statistical power for investigating the differences between pipelines, the results across datasets are aggregated to conduct a meta-analysis. Our Phase-SPDNet yields statistically better results than all other approaches as shown in Fig. 3 (b), where only significant interactions are displayed ($p$ > 0.0.5), the reported values and color show the $t$-values.
@@
The second relevant point is to see the impact of the augmentation procedure compared to the standard SPDNet. Figure 3 (a) elucidates this aspect by illustrating the relative performance enhancements of various models against the SPDNet standard. Notably, ShallowNet and DeepNet exhibit positive performance increments across the majority of the datasets examined. However, the augmentation procedure, as implemented by Phase-SPDNet , not only consistently yields positive relative improvement against the standard SPDNet across all datasets but also the improvement introduced by the augmentation procedure turns out to be so significant that it even outperforms methods that initially surpassed the SPDNet benchmark. This improvement is not only demonstrated by the relative improvement with respect to the SPDNet standard but also is evident through the meta-analysis comparison (refer to Fig 3 (c), (d) and (e)).
@@
Ultimately, the statistical analysis conducted using Figure 3 (b) allows us to demonstrate that Phase-SPDNet outperforms the state-of-the-art DL methodology in a reduced dataset context. Its versatility and reliability, particularly in data-limited scenarios, position the method as a promising solution for real-world applications.
@@
Table 2: Average AUC-ROC (%) on six datasets (BCNI2014001, BCNI2014004, Cho2017, Schirrmeister2017, Weibo2014, and Zhou2016), for a left vs. right motor imagery task. The highest performances are in bold-face.
@@
Phase-SPDNet (Our) 75.98 $\pm$ 17.00 80.46 $\pm$ 16.64 66.00 $\pm$ 13.29 72.02 $\pm$ 13.07 78.01 $\pm$ 19.64 94.92 $\pm$ 3.34
Figure 3: Results for Right vs Left-hand classification, using Within-Session evaluation. Plot (a) provides the relative improvement of the AUC-ROC in percentage of the method considered with respect to the standard SPDNet of the different pipelines considered. Plot (b) shows a combined meta-analysis (over all datasets) of the different pipelines. It shows the significance of the algorithm on the y-axis being better than the one on the x-axis. The gray level represents the significance level of the ROC-AUC difference in terms of t-values. We only show significant interactions ($p<0.05$). Plots (c), (d), and (e) show the meta-analysis of Phase-SPDNet against SPDNet, ShallowNet, and DynSpat+ShallowNet, respectively. We show the standardized mean differences of p-values computed as a one-tailed Wilcoxon signed-rank test for the hypothesis given in the plot title. The gray bar denotes the $95\%$ interval. * stands for $p<0.05$, ** for $p<0.01$, and *** for $p<0.001$.
@@ 5.1 Impact of augmentation on most and least responsive 5 subject of Cho2017
In the previous section, we explored the outcome obtained by Phase-SPDNet within the context of a reduced number of electrodes. Now, our attention shifts to a deeper exploration of this new algorithm, with a specific emphasis on the interpretability and computational performance of the method. To facilitate this comprehensive analysis, still in the context of three electrodes, we have chosen to concentrate on the examination of the five most and least responsive subjects within the Cho2017 dataset (the wider dataset in terms of the number of subjects), selected using the Minimum Distance to the Mean (MDM) algorithm [63] with covariance as a feature.
To analyze the performance, we use different pipelines, MDM [63], ACM+MDM [7], SPDNet, and Phase-SPDNet , each offering a distinct perspective on the impact of augmentation techniques on both traditional ML and advanced DL algorithms. Furthermore, we assess the performance using three different SPD features: covariance, instantaneous coherence, and imaginary coherence (Figure 4).
Figure 4: (a) Plot showing the box plot of the five most responsive subjects (3, 14, 35, 41, 43) of Cho2017. We see that the augmentation procedure consistently leads to notable improvements in the classification performance for all the SPD estimators considered.(b) Plot showing the box plot of the least responsive five subjects (2, 7, 29, 34, 50) of Cho2017. We see that the augmentation procedure increases the performance for the covariance feature, while the improvement is less clear for the FC estimator. In the plot, the names Ins and stand respectively for Instantaneous Coherence and Imaginary Coherence.
For the five most responsive subjects, covariance stands out as the top-performing feature as it could be expected. However, in the case of the five least responsive subjects, this dominance of covariance is less clear when considering the state-of-the-art methodology.
@@
It becomes evident that the augmentation procedure consistently leads to notable improvements in the classification performance for the top-performing five subjects. However, the situation becomes less definitive for the least responsive five subjects, where the augmentation effect is not as pronounced or evident. In this case, their signals exhibit a lower SNR, which can substantially impact the MDOP algorithm, leading to a poor estimation of the hyper-parameter. In this case, the augmentation procedure achieves a very positive impact on covariance while there is even a decrease in performance using coherence.
@@ 5.2 Interpretability
Creating explainable algorithms is crucial in the field of DL, and it holds even greater significance in healthcare data. In healthcare, understanding why an algorithm makes specific decisions is of great importance for clinicians and patients, ensuring transparency, accountability, bias mitigation, education, and trust-building in the collaboration between AI systems and medical experts.
For analyzing the explainability of the Phase-SPDNet we use the GradCam++ algorithm [86], a technique designed for visualizing and explaining the decision-making process of neural networks. The GradCam++ output is essentially a visual representation that reveals which elements within the data are the most important for the model. GradCam++ has proven to be a robust method for spatial interpretability of EEG signals in a class-agnostic manner [87].
@@
In our analysis, we focus on Subject 41 from the Cho2017 dataset, which is notable for being among the top 5 subjects that exhibit the most substantial performance improvement through the augmentation procedure. Similar findings also hold for other subjects. In particular, using covariance as a feature, we show a sample of the test dataset that was correctly classified by the Phase-SPDNet algorithm (left hand classified as left hand) but was misclassified by SPDNet using the same feature, as shown in Figure 5. Similar findings have been observed for the FC estimators. This figure represents the output of the GradCam++ applied to the output of the ReEig layer.
@@
In all three scenarios, the standard SPDNet places more emphasis on the diagonal terms of SPD matrices. On the other side, when employing Phase-SPDNet , a more comprehensive picture unfolds. While diagonal terms continue to play a pivotal role, what becomes increasingly evident is the significant contribution of the off-diagonal terms. In essence, this observation emphasizes that the Phase-SPDNet model leverages both individual electrode characteristics (diagonal terms) and the interplay between electrodes in time and space (off-diagonal terms) as crucial elements in its classification decision-making process.
@@
Figure 5: Interpretability parallel between GradCam++ and t-test between right hand left-hand classification for subject 41 of Cho2017. Plot (a) is the GradCam++ visualization, applied after the ReEig layer, for Augmented covariance. Plot (b) is a zoom in the region corresponding to standard Covariance. Plot (c) is the GradCam++ visualization, applied after the ReEig layer for standard covariance. Plot (d) is the t-test of the Augmented covariance matrix. Plot (e) is the t-test of the covariance matrix. The t-test visualization represents only the significant interactions ($p<0.05$). For the standard covariance matrix, the only significant interaction occurs between electrodes C3, which mirrors the representation seen in the GradCam++ visualizations. In contrast, the Phase-SPDNet (Figure 5(d)) reveals additional significant interactions, although this interaction map does not precisely match the GradCam++ visualization. This discrepancy highlights the added value of our method in identifying crucial inter-channel relationships that might be overlooked by traditional covariance analysis.
To investigate deeper into the interpretability of our proposed network, we employ a paired t-test to highlight statistically significant differences in the augmented covariance matrix (ACM) between left and right hands. Our analysis centers on subject 41 from Cho2017, specifically on the test dataset of the initial fold of the cross-validation, with consistent findings observed in the other folds. Given that 20 samples are available for each imagined movement, we are in the domain of application of the t-test.
Examining Figure 5 (d) and (e), it is evident that SPDNet places emphasis primarily on the statistically significant pair of electrodes, showing a clear parallel between the GradCam++ visualization and the t-test. In contrast, Phase-SPDNet not only considers significant interactions but also extends its focus to nearly significant interactions.
@@ 5.3 Convergence Behavior
To analyze convergence behavior, we again consider subject 41 of the Cho2017 dataset. Again, to get a fair comparison, all the following results are calculated on the same hardware, a Dell C6420 dual-Xeon Cascade Lake SP Gold 6240 @ 2.60GHz.
We examine the outcomes concerning both training and validation metrics, including loss and Receiver Operating Characteristic Area Under the Curve (ROC AUC) scores. The consolidated results are visually presented in Figure 6, wherein the graphs depict the average values across the considered 5-fold cross-validation.
Notably, our observations reveal that the loss function of Phase-SPDNet converges to zero at a notably accelerated pace and exhibits a considerably more stable trajectory compared to standard SPDNet. Moreover, while the loss function of Phase-SPDNet steadily approaches zero, the standard SPDNet experiences a more gradual descent, deviating from the path to complete convergence but ultimately plateauing around $0.4$. These findings highlight that the capacity of SPDNet is limited and that the algorithm stops learning at a certain point.
Interestingly, when examining the ROC AUC graph, we observe that the validation scores of Phase-SPDNet converge consistently to higher values. Notably, this convergence is achieved even faster compared to the SPDNet standard, requiring only $94\%$ of the iterations. This implies that Phase-SPDNet attains a comparable performance to SPDNet but with a significantly reduced number of iterations, making it an efficient and effective choice for real-life applications.
Figure 6: Convergence behavior for right-hand vs. left-hand classification using WS evaluation. The plot uses subject 41 of the Cho2017 dataset, showing the mean value for each plot computed over the 5-Fold considered in the WS evaluation. Plot (a) provides the training loss, while plot (b) is the validation one. Plot (c) provides the ROC-AUC for the training phase, while plot (d) the one for the validation.
5.4 Computational Performances
In this section, we are interested in the analysis of computational performance, with a specific focus on two key aspects: execution time and environmental impact. The results on the best 5 subjects of Cho2017 (subjects 3, 14, 35, 41, 43) are presented in Table 5.4.
Recently, much attention has begun to be given to the sustainability of ML and AI algorithms [88]. Since more and more models continue to proliferate and demand substantial computational resources, it is crucial to estimate and mitigate the energy consumption and emissions associated with their training and deployment. To assess the environmental impact of our algorithm, we use the CodeCarbon [89] library. CodeCarbon is an open-source library that enables developers to monitor carbon dioxide ($CO_{2}$) emissions, offering a way to analyze the carbon footprint of our work.
Interestingly, our analysis shows that the Phase-SPDNet exhibits the longer computational time among the evaluated algorithms while simultaneously standing out as one of the most environmentally friendly in terms of its $CO_{2}$ impact (see Fig 5.4). This intriguing juxtaposition can be attributed to the fact that, during each fold of the cross-validation process, there’s a necessity to optimize the model’s hyper-parameters, which naturally augments the computational time. However, the MDOP algorithm consumes fewer resources than gradient descent procedures, thus resulting in a comparatively smaller environmental impact. In fact, the use of augmented procedure introduces the complication of two hyper-parameters but nevertheless manages to keep the number of trainable parameters of the model contained (see Table 5.4). It’s plausible that our algorithm has yet to reach its optimal performance. Further investigation is needed to improve the optimization strategies.

### v3

@@ 3.6 Datasets
We employed standard pre-processing steps designed for all datasets described in [10]. The pre-processing steps are the same for all the pipelines considered in our study. We applied band-pass filtering between $[8-32]$ Hz [83, 84] and electrode standardization, bringing every channel to a zero mean and a unit standard deviation. We opted to use the complete epoch duration, although this duration varies among datasets, as shown in Table 1.
We concentrated on a binary classification task, aiming to discern between imagined movements of the right and left hand using an intra-subject/within-session evaluation, which is based on a $5-$fold cross-validation conducted for each session.
We also focus on investigating our algorithm’s robustness with fewer electrode settings. For this purpose, our channel selection process was guided by two constraints: ensuring neurophysiological relevance and maintaining consistent electrode selection across datasets. We selected electrodes positioned over the sensorimotor region, particularly in the central (C) and centro-parietal (CP) lines, which are actively involved in the tasks. This choice is strongly supported by BCI research [84, 85, 86, 87, 88, 89, 90]. In the considered datasets we used EEG montages with varying numbers of electrodes, from 3 to 128. We standardized our analysis by using the minimal overlapping electrodes common to all dataset: C3, Cz, and C4. All datasets considered were based on reference electrodes, and no average reference was used.
@@ 3.7 Baseline Comparison
These methods were implemented and trained using MOABB version 1.1.0 [10]. For in-depth information about the deep learning hyper-parameters, please consult Table A2.
@@ 4 Results
In this section, we analyze the results produced using within-session (WS) evaluation. This method is based on a 5-fold cross-validation performed on each session independently. In particular, we report two separate analyses.
@@
The first analysis is dedicated to contrasting the performance of Phase-SPDNet , showing both MDOP and optimized search, with a specific focus on the exploitation of the covariance feature, against the state-of-the-art techniques in DL applied in EEG. The subsequent analysis, reported in the Appendix, shifts the focus towards the exploration of coherence as a feature for classification. Despite the ongoing investigations into coherence, we deliberately separated this analysis due to its comparatively lesser performance against the benchmarks achieved with covariance features. Nevertheless, the augmentation procedure is a significative methodology to improve the performance of both feature sets.
@@
For the first analysis, we tested the performance of Phase-SPDNet MDOP and Phase-SPDNet OPT against the baseline comparison defined in 3.7 across a broad spectrum of datasets. The results are listed in Table 2, while the detailed statistical analysis can be found in Figure 3.
Table 2 shows ROC AUC results for right-hand versus left-hand classification, where bold numbers represent the best score in each dataset, revealing a consistent trend wherein our algorithm Phase-SPDNet OPT surpasses the DL state-of-the-art in the context of reduced datasets in all considered datasets. The statistical analysis conducted, shown in Figure 3 (b), allows us to demonstrate that Phase-SPDNet OPT outperforms the state-of-the-art DL methodology in a reduced dataset context.
When employing MDOP, Phase-SPDNet MDOP obtains suboptimal performance with respect to Phase-SPDNet OPT, finding aligned with current research [7], while still surpassing the performance of the DL state-of-the-art. A notable exception is the Schirrmeister2017 dataset, where Phase-SPDNet MDOP is marginally surpassed by ShallowNet, DeepNet, and DynSpat-ShallowNet. However, we are interested in computing a statistical behavior considering all datasets under consideration. When considering the aggregated results across all datasets, our meta-analysis shows that Phase-SPDNet MDOP yields statistically better results than all other approaches, as shown in Fig. 3 (b).
@@
The second relevant point is to see the impact of the augmentation procedure compared to the standard SPDNet. Figure 3 (a) elucidates this aspect by illustrating the relative performance enhancements of various models against the SPDNet standard. Notably, ShallowNet and DeepNet exhibit positive performance increments across the majority of the datasets examined. However, the augmentation procedure, as implemented by Phase-SPDNet OPT, not only consistently yields positive relative improvement against the standard SPDNet across all datasets but also the improvement introduced by the augmentation procedure turns out to be so significant that it even outperforms methods that initially surpassed the SPDNet benchmark. This improvement is not only demonstrated by the relative improvement with respect to the SPDNet standard but also is evident through the meta-analysis comparison (refer to Fig 3 (c), (d) and (e)).
@@
Ultimately, the versatility and reliability of Phase-SPDNet OPT, particularly in data-limited scenarios, position the method as a promising solution for real-world applications.
@@
Table 2: Average AUC-ROC (%) on six datasets (BCNI2014001, BCNI2014004, Cho2017, Schirrmeister2017, Weibo2014, and Zhou2016), for a left vs. right motor imagery task. The highest performances are in bold-face. The gray area marks the second to best pipeline excluding the Phase-SPDNet MDOP.
@@
Phase-SPDNet MDOP (Our) 75.98 $\pm$ 17.00 80.46 $\pm$ 16.64 66.00 $\pm$ 13.29 72.02 $\pm$ 13.07 78.01 $\pm$ 19.64 94.92 $\pm$ 3.34
Phase-SPDNet OPT (Our) 78.40 $\pm$ 15.12 82.29 $\pm$ 17.04 66.69 $\pm$ 12.99 76.05 $\pm$ 13.84 78.18 $\pm$ 19.36 95.62 $\pm$ 2.36
Figure 3: Results for Right vs Left-hand classification, using Within-Session evaluation. Plot (a) provides the relative improvement of the AUC-ROC in percentage of the method considered with respect to the standard SPDNet of the different pipelines considered. Plot (b) shows a combined meta-analysis (over all datasets) of the different pipelines. It shows the significance of the algorithm on the y-axis being better than the one on the x-axis. The gray level represents the significance level of the ROC-AUC difference in terms of t-values. We only show significant interactions ($p<0.05$). Plots (c), (d), and (e) show the meta-analysis of Phase-SPDNet OPT against SPDNet, ShallowNet, and DynSpat+ShallowNet, respectively. We show the standardized mean differences of p-values computed as a one-tailed Wilcoxon signed-rank test for the hypothesis given in the plot title. The gray bar denotes the $95\%$ interval. * stands for $p<0.05$, ** for $p<0.01$, and *** for $p<0.001$.
@@ 5.1 Impact of augmentation on most and least responsive 5 subject of Cho2017
In the previous section, we explored the outcome obtained by Phase-SPDNet within the context of a reduced number of electrodes. Now, our attention shifts to a deeper exploration of this new algorithm. Still, in the context of three electrodes, we have chosen to concentrate on the examination of the five most and least responsive subjects within the Cho2017 dataset (the wider dataset in terms of the number of subjects). These subjects were selected on the basis of the MDM algorithm [69] with covariance as a feature.
To analyze the performance, we use different pipelines, MDM [69], ACM+MDM [7], SPDNet, and Phase-SPDNet OPT, each offering a distinct perspective on the impact of augmentation techniques on both traditional ML and advanced DL algorithms. Furthermore, we assess the performance using three different SPD features: covariance, instantaneous coherence, and imaginary coherence (Figure 4).
Figure 4: Most and least responsive 5 subjects of Cho2017. (a) Plot showing the box plot of the five most responsive subjects (3, 14, 35, 41, 43) of Cho2017. We see that the augmentation procedure consistently leads to notable improvements in the classification performance for all the SPD estimators considered. (b) Plot showing the box plot of the least responsive five subjects (2, 7, 29, 34, 50) of Cho2017. We see that the augmentation procedure increases the performance for the covariance feature, while the improvement is less clear for some FC estimators. In the plot, the names Ins and Imn stand respectively for Instantaneous Coherence and Imaginary Coherence.
For the five most responsive subjects, covariance stands out as the top-performing feature, as could be expected. However, in the case of the five least responsive subjects, this dominance of covariance is less clear when considering the state-of-the-art methodology. Nevertheless, considering the Phase-SPDNet OPT method that uses covariance as a feature, we see that covariance is again the best-performing feature.
@@
The augmentation procedure consistently leads to notable improvements in the classification performance for the top-performing five subjects. However, the situation becomes less definitive for the least responsive five subjects, where the augmentation effect is evident for the covariance but is not as pronounced for the FC estimator. In this case, their signals exhibit a lower SNR, which can substantially impact the MDOP algorithm, leading to a poor estimation of the hyper-parameter.
@@ 5.2 Interpretability
For analyzing the explainability of Phase-SPDNet , focusing on the best performing Phase-SPDNet OPT, we use the GradCam++ algorithm [92], a technique designed for visualizing and explaining the decision-making process of neural networks. The GradCam++ uses the derivative of the gradient for a specific target (e.g., right hand or left hand) that flows up to a specific convolutional layer. This process creates a map highlighting the critical regions used to predict that specific target class, without requiring architectural modifications or re-training. GradCam++ has proven to be a robust method for spatial interpretability of EEG signals in a class-agnostic manner [93].
@@
In our analysis, we focus on Subject 35 from the Cho2017 dataset, which is notable for being among the top 5 subjects that exhibit the most substantial performance improvement through the augmentation procedure. Similar findings also hold for other subjects. We create the GradCam++ map, produced with left hand as a target, on a sample of the test dataset that was correctly classified by the Phase-SPDNet OPT algorithm (left hand classified as left hand) but was misclassified by SPDNet using the same feature, as shown in Figure 5. The obtained map represents the GradCam++ obtained at the ReEig layer.
@@
Standard SPDNet places more emphasis on the diagonal terms of SPD matrices. On the other side, when employing Phase-SPDNet OPT, a more comprehensive picture unfolds. While diagonal terms continue to play a pivotal role, what becomes increasingly evident is the significant contribution of some off-diagonal terms. Moreover, the Phase-SPDNet with optuna emphasizes the interplay between a channel and its lagged version, showing that this information is even more relevant for the classification with respect to the standard covariance itself.
In essence, this observation emphasizes that the Phase-SPDNet OPT model leverages both individual electrode characteristics (diagonal terms) and the interplay between electrodes in time and space (off-diagonal terms) as crucial elements in its classification decision-making process.
@@
Figure 5: Interpretability parallel between GradCam++ and t-test for right versus left-hand classification for subject 35 of Cho2017. Plot (a) is the GradCam++ saliency map obtained at the level of the ReEig layer, for Phase-SPDNet OPT with a target left hand. Plot (b) is a zoom in the region corresponding to standard Covariance. Plot (c) the GradCam++ saliency map obtained at the level of the ReEig layer, for SPDNet with a target left hand. This discrepancy highlights the added value of our method in identifying crucial inter-channel relationships that might be overlooked by traditional covariance analysis.
@@ 5.3 Convergence Behavior
To analyze convergence behavior, we again consider subject 35 of the Cho2017 dataset. We examine the outcomes concerning both training and validation loss. The consolidated results are visually presented in Figure 6, wherein the graphs depict the average values across the considered 5-fold cross-validation.
Notably, our observations reveal that the loss function of Phase-SPDNet OPT shows a rapid decline in the loss function during the initial iteration, suggesting it learns more quickly than the SPDNet model. Moreover, Phase-SPDNet OPT steadily decreases the loss, while the SPDNet decreases initially but plateaus around a higher value compared to the proposed approach. These findings highlight that Phase-SPDNet OPT possesses a better learning efficiency and generalization capability.
Figure 6: Convergence behavior for right-hand vs. left-hand classification using WS evaluation. The plot uses subject 35 of the Cho2017 dataset, showing the mean value for each plot computed over the 5-Fold considered in the WS evaluation. Plot (a) provides the training loss, while plot (b) is the validation one.
5.4 Trainable Parameters
The proposed methodology has demonstrated a statistically significant enhancement in performance without compromising efficiency. Indeed, this model features a reduced number of trainable parameters, particularly influenced by the hyperparameter $\psi$, as shown in Table 5.4. The table offers an in-depth analysis of how variations in $\psi$ affect the overall architecture.
While there is an observable increase in the number of trainable parameters when compared to the conventional SPDNet, this increase is balanced by a significant improvement in performance. Despite the additional parameters relative to the SPDNet standard, our model remains considerably more efficient in terms of parametric complexity than the current state-of-the-art DL pipelines, as shown in Figure 5.4.